HighJump Inc. is considering partnering with FieldFeet to open "Athlete Heaven" stores in the United States as it seeks to expand its own retail network to better control how its products are displayed and sold. One particular store will require an initial investment of $220,000 and an annual operating cost of $59,000. The buildings and equipment will have a $62,500 resale value after 10 years. HighJump expects the store to generate annual revenue of $69,000. Calculate the future worth of the investment in this particular store at the end of year 10, and determine the acceptability of the investment if the company's minimum attractive rate of return is 8% per year.
Answer #1

i = 8%

Present worth of the store = -220000 + (69000 - 59000)*(P/A,8%,10) + 62500*(P/F,8%,10)

= -220000 + 10000*(P/A,8%,10) + 62500*(P/F,8%,10)

= -220000 + 10000*6.710081 + 62500*0.463193

= -123949.59

Future worth after 10 years = -123949.59 * (F/P, 8%,10)

= -123949.59 * 2.158924

= -267597.87

As future worth is negative, therefore project is not acceptable